OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). 2 4 −2 −8

OpenStudy (anonymous):

( F+g)(-1) = f(-1) + g(-1)

OpenStudy (anonymous):

f(-1) = 2 * -1 +1 g(-1) = -2 *-1 -3 Add =-2

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

If u r Satisfied give medals by clicking best answer option